OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@sleepyjess plz come help

OpenStudy (sleepyjess):

as i said last night, we need to find the slope first do you know how to do that?

OpenStudy (anonymous):

not reallly

OpenStudy (sleepyjess):

you have 6, 2 as point 1 and 4, 8 as point 2 \(\dfrac{y_2-y_1}{x_2-x_1}\) 6, 2 = \(x_1,~y_1\) 4, 8 = \(x_2,~y_2\)

OpenStudy (anonymous):

then wat happens after that to find the answer

OpenStudy (sleepyjess):

substitute the points i gave you into the formula

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@sleepyjess i didnt get the answer i dont get hoiw to substitute it

OpenStudy (sleepyjess):

show me what you did to substitute the numbers i gave you the points and what the point represented

OpenStudy (anonymous):

but how do i susbitute

OpenStudy (sleepyjess):

use 6 as \(x_1\) 2 as \(y_1\) 4 as \(x_2\) and 8 as \(y_2\)

OpenStudy (anonymous):

@sleepyjess it has to be like this y = x +

OpenStudy (sleepyjess):

i know but to get that we need to get the slope

OpenStudy (anonymous):

so wat would be the sope plz

OpenStudy (sleepyjess):

just put in the numbers to the formula i gave you

OpenStudy (anonymous):

sooo 6,2 is 1x i think

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@sleepyjess is this it

OpenStudy (sleepyjess):

\(\dfrac{8-2}{4-6}\)

OpenStudy (anonymous):

so 6x -2

OpenStudy (anonymous):

@sleepyjess so 6x -2

OpenStudy (sleepyjess):

no your slope is \(-\dfrac{6}{2}\) which can be simplified to -3 now we have -3x+? can you find the y intercept with that information?

OpenStudy (anonymous):

um 16

OpenStudy (anonymous):

@sleepyjess is is 16 or -16

OpenStudy (sleepyjess):

going down 2 for x and up 6 for y point 3 is 2, 14 point 4 is 0, 20 which point is the intercept?

OpenStudy (anonymous):

it says im typin bt i9m not

OpenStudy (sleepyjess):

no which point has the x value as 0? 3 or 4?

OpenStudy (sleepyjess):

its fine it does it for me too

OpenStudy (anonymous):

ummm 8

OpenStudy (anonymous):

o 4

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

i said 4 not 8

OpenStudy (sleepyjess):

yes point 4 has an x value of 0 what is the y value for point 4?

OpenStudy (anonymous):

@sleepyjess u still here

OpenStudy (anonymous):

um 20

OpenStudy (sleepyjess):

yes so your equation is y=-3x+20

OpenStudy (anonymous):

20

OpenStudy (anonymous):

@sleethx i understand that now i got diff. kind of quesion kjind of diff.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 1 −2 −5 −8 y = x +

OpenStudy (anonymous):

@sleepyjess srry abought that lag plz come help me

OpenStudy (sleepyjess):

sorry my computer decided to randomly restart

OpenStudy (testing123guest):

hay man can u help me out with that first question like what was the answer it was not that clear